a rectangular water tank has dimensions of 2.0 m x 1.0 m x 1.0 m. what would be the diameter of a 1.50-m high cylinder tank to contain the same amount of water that the rectangular tank contains?
1
Expert's answer
2020-09-29T09:39:33-0400

The volume of water, that fits into rectangular tank is "V_r = 2 m \cdot 1 m \cdot 1 m = 2 m^3".

Volume of cylinder is "V_c = \pi \left(\frac{d}{2}\right)^2 H", where "d" is diameter and "H"is height.

Hence, given volume of cylinder and height, the diameter must be "d = 2 \sqrt{\frac{V_c}{\pi H}}". Replacing the volume of cylinder with the volume of rectangular tank, obtain: "d = 2 \sqrt{\frac{V_r}{\pi H}} \approx 1.3m".